CERN/LEPC 99-6
LEPC 52
9 September 1999

 
 


LEP EXPERIMENTS COMMITTEE

Minutes of the 52nd meeting,
7 September 1999

OPEN SESSION

1. Status of LEP operation P. Collier reported on the status and performance of LEP. This year LEP has operated at beam energies of 96 GeV (for a total luminosity of 31 pb-1 per experiment) and 98 GeV (86 pb-1), and is currently at 100 GeV (50 pb-1, at the time of the presentation). After including the Z calibration data (3.2 pb-1), the total delivered luminosity so far was 171 pb-1. Not only was the record energy of 100 GeV per beam achieved more than one month ahead of schedule, but records were broken for maximum beam-beam tune shift (0.083), beam current (6.6 mA) and luminosity (1.2x1032 cm-2s-1, 4 pb-1 per day and 19 pb-1 per week). These performances have been achieved with the 102/90 optics, which works well for both Z calibration data and high energy data.

An rf accelerating voltage of 3420 MV is now available at LEP from the full system of 272 Nb/Cu sc cavities (3205 MV), 16 Nb sc cavities (104 MV) and 48 warm Cu cavities (110 MV). The Nb/Cu cavities are currently operating at 7 MV/m mean gradient - well above their design value of 6 MV/m. With this rf voltage, LEP should be able to reach 101 GeV per beam during the last few weeks of the scheduled running time, assuming INB authorisation to exceed 100 GeV is received from the French authorities.

Despite the excellent overall performance, some difficulties have been experienced. The new cryogenics system, upgraded this year from 6 to 12 kW, has had several problems, including water in the He circuit which necessitates weekly de-iceing of the filters. Six of the 44 klystrons have suffered failures. Three breakdowns have occurred in 18 kV transformers, causing major power cuts and beam losses totalling 3.5 days. These difficulties may reflect the fact that LEP is being pushed to its limits.

During the 99/00 shutdown several improvements will be made to allow a further increase of beam energy in 2000. These include installation of 8 more Cu cavities, equipping the horizontal correctors with power converters so they act as ring dipoles, processing of some sc rf modules and modifications to the cryogenics system. Work is also continuing on the development of a new low emittance lattice (131/90 optics) which could potentially provide higher beam energies.

2. Status of beam energy calibration with polarisation and spectrometer

B. Dehning reported on the status of the LEP2 beam energy calibration with the precision spectrometer. The method involves first calibrating the spectrometer at a known beam energy using resonant beam depolarisation and then using this calibration to precisely measure the deflection angle of the beam at higher energy. To measure the beam energy with a precision s < 10 MeV, the dipole magnet field integral must be measured to a precision 3x10-5 at each of its operating values, and the beam position monitor (BPM) resolution and stability must be below 1 mm. Both these figures have been achieved, although present variations of the BPM gain with beam current have to be understood.

3. LEP detector status reports
 

 L3:           G. Bobbink
OPAL:       M. Hauschild
ALEPH:    O. Callot
DELPHI :  P. Charpentier


The four experiments reported good overall background conditions at 200 GeV and high data collection efficiencies of 90-95%, which have resulted in data samples of about 150 pb-1 so far this year. All detectors enthusiastically support the operation of LEP in the two-week reserve period at ÷s = 202 GeV, which will provide about an extra 2 GeV in mass sensitivities for Higgs and SUSY particles.

Three experiments reported damage to their central tracking detectors. OPAL suffered a spark before LEP startup which has resulted in permanent field distortions. However, these are now fully corrected. Both ALEPH and L3 reported serious damage to their central tracking chambers - in some cases correlated with beam losses - resulting in permanent HV breakdown in specific regions of the chambers. Two such incidents have occurred in L3 (although the second did not involve significant beam loss), necessitating a permanent voltage reduction (80-90%) in 4 of the 24 main sectors of the TEC. This has severely affected the TEC calibration, requiring further Z data. ALEPH plans a major detector access during the coming winter shutdown to repair the TPC short.

Concerning physics results after LEP shutdown, most experiments expect to complete their full analysis by the end of 2003, although OPAL foresees analysis continuing until 2006. The four detectors request CERN to support the current analysis tools (hardware and software) during this period, even though this may involve some difficulties with obsolete equipment, and also to provide appropriate support for the experimental research groups.
 

4. LEP working group reports

Higgs:   P. McNamara

SUSY:   G. Ganis
 

Reports were presented on the combined data from the present run by the two "search" working groups as a test of the fast combination procedures, in preparation for the final year of LEP operation. The combined analyses were successfully carried out in a period of about 3 weeks, which can probably be reduced for next year. No signs of new physics have been observed. Using data up to ÷s = 196 GeV, the combined lower mass limit on the SM Higgs is 102.6 GeV (95% CL). The lower mass limits on the lightest Higgs of minimal SUSY (h,A) are near 84.5 GeV for large values of tan b; the low tan b range of 0.5-3.2 is excluded for no-mixing scenarios. The charged Higgs is excluded below 77 GeV.

5. CosmoLEP

K. Eggert presented the CosmoLEP proposal. The primary physics aims of the experiment are a) to measure the proton vs. iron ratio at the knee of the cosmic ray energy spectrum (near 3x1015 eV), and b) to look for surprises. The former question is of interest since it may help explain the origin of cosmic rays. The proposed technique is to study multi-muon bundles with muon energies above about 70 GeV, corresponding to the energy cutoff at a depth of 140 m. The experiment involves the installation of a large area (160 m2) muon chamber array in the cavern adjacent to ALEPH. The transverse profile and multiplicity of the multi-muon events will be measured, both with and without a coincidence with ALEPH

.
CLOSED SESSION

Present: R. Cashmore, J. Colas, M. Delfino, F. Gasparini, V.G. Goggi, N. Harnew,
K. Hübner, P.O. Hulth, P. Janot, L. Jonsson, J. Kirkby (Secretary), K.-H. Kissler,
W. Lohmann, L. Maiani, M. Mangano, C. Matteuzzi, J. May, R. Miquel, T Müller, S. Myers,
Y. Sirois, A. Smith, M. Spiro (Chairman), R. Tenchini, A. Watson

Apologies: M. Doser, D. Schaile
 

1. Discussion on the CosmoLEP proposal

The Chairman was pleased to announce that two cosmic ray experts - Tom Gaisser and Alan Watson - had agreed to referee the CosmoLEP proposal for the LEPC. Although both were presently attending the TAUP conference in Paris, A. Watson had been able to make a special trip to be present for the presentation and discussion of CosmoLEP, and to talk with the proponents. In view of his time constraint, the discussion of the proposal was brought forward on the agenda.

In the course of a lively discussion, several open questions were raised by the committee. In particular, does the proposed experiment have the sensitivity to solve the question of the p-Fe ratio at the knee given the limited experimental quantities that would be measured and given the uncertainties in the Monte Carlo generators of energetic cosmic ray showers in the atmosphere? The committee is also interested to hear the special advantages of CosmoLEP compared with current cosmic ray experiments that are presently addressing the same physics question. It was agreed that the referees would discuss these and other questions with the proponents and make an oral report to the committee at the November meeting, where a final recommendation would be made.

2. Approval of the minutes of the 51st meeting

The minutes of the 51st meeting (LEPC 99-3/LEPC 51) were approved without modification.

3. Chairman's report

The Chairman was pleased to announce that Robert Pittau has agreed to chair the LEP200 Monte Carlo working group.
 

4. Discussion on the LEP machine reports

The Chairman joined with the committee in warmly congratulating the machine team for the record-breaking performance of LEP this year. Concerning the LEP2 beam energy calibration, the committee appreciated the good progress on the precision beam spectrometer. The committee was also pleased to hear the new technique to determine the LEP2 beam energy from the frequency of the longitudinal oscillations of the beam (Qs) and the total rf voltage. This method has already produced an independent measurement of the LEP2 beam energy with an uncertainty comparable to the flux loop and NMR methods (s = 20 MeV).
 

5. Discussion on the LEP detector status reports

The committee noted the excellent operational efficiency of the LEP detectors at the new peak energies and was impressed by the high data quality, which allowed fast physics results to be obtained by the working groups. However the committee was disturbed by the damage incurred by two of the detectors (ALEPH and L3) due, in part, to beam losses. No new procedures could be identified from the machine side that could have led to this exceptionally high incidence of beam-loss damage. The committee acknowledged the serious impact this has had on the calibration of the tracking detectors for ALEPH and L3, and supports the LEP coordinator to decide, in consultation with the four collaborations, on any additional Z calibration data that may be required before the end of the run beyond the 0.5 pb-1 presently scheduled.

Concerning completion of the analysis of LEP data after machine shutdown, the committee noted that all four experiments wish to maintain their full analysis system in the period 2001-2003. After this period some experiments foresee further full analysis and some would convert to archive-only analysis. The committee encourages CERN to maintain the current analysis platforms of the LEP experiments during the period 2001-2003, while recognising this may involve some outdated equipment and software, and also to provide appropriate support for the experimental research groups over the same period.
 

6. Discussion on the LEP working group reports

The committee was pleased to hear the success of the search working groups in producing fast combined results from LEP data taken this year. These analyses were impressively completed in a period of only about 3 weeks since the LP99 meeting in August.

A strategy has been agreed among the experiments for the combined searches next year. In the case of the Higgs this will involve essentially a continual combined assessment by the working group of the few channels involved. In the case of SUSY, where many channels are involved, it has been agreed that any 2-3s signal appearing in an individual experiment will be presented within the working group and trigger a combined search in the four experiments. The committee fully supports this strategy for the search working groups.
 

7. Report from the LEP Coordinator

The LEP Coordinator reported that there have been 108 days so far for physics this year, with about 15 days lost for various reasons. A total of 39 days remain for physics, plus 14 days in reserve. The total delivered luminosity is presently 171 pb-1 per experiment and the variation between experiments is now excellent (within ±2%). On the other hand, there are worries concerning transformer-induced power cuts (which result in electronics failures), damage in the detectors due partly to beam losses, and cryogenics failures (affecting the ALEPH and DELPHI magnets).

Concerning the 2-week LEP reserve, the machine team and all four detectors support the proposal to operate at a beam energy of 101 GeV, assuming the INB authorisation is received. Excluding data at 101 GeV, by the end of the present cycle the expected SM Higgs combined 5s discovery mass limit will be 102.5 GeV and combined exclusion limit will be 106.5 GeV at 95% CL. Assuming 2 weeks at 101 GeV providing about 15 pb-1 (without the 2-week reserve), the limits will be 103.5 GeV and 107.5 GeV, respectively. Adding the 2-week reserve at 101 GeV would raise these limits to 105 GeV and 109 GeV, respectively, indicating significant new discovery potential.

Several additional arguments were raised in support of running LEP at 101 GeV during the 2-week reserve. They included the expectation that the available rf voltage would likely be increased during the scheduled stop in week 41 (which should improve the machine efficiency), the added experience with the machine prior to the winter shutdown, during which any unforeseen problems could be remedied, the current losses of about 2 weeks so far this year with the consequent financial savings and, not least, the excellent current state of operation of LEP. The committee agreed that a compelling case exists and strongly supports the operation of LEP during the 2-week reserve at 101 GeV.
 

8. Next LEPC meeting

The dates of the next meeting of the LEPC are Tuesday-Wednesday, 9-10 November.

The tentative dates for the LEPC in 2000 are
 

07 March (physics jamboree; revised dates)

20 July (combined search results)

either 16 - 18 October (LEP symposium)
or 13 - 15 November (LEP symposium)


The special meeting on 20 July will include presentations of the combined data from the four experiments in order to allow a timely recommendation for any extension of LEP beyond the planned shutdown date in September 2000. The final LEPC meeting in 2000 will coincide with a 3-day symposium planned to celebrate LEP (for which the dates are not yet finalised).
 

J. Kirkby
DOCUMENTS RECEIVED [1] C. Taylor et al., CosmoLEP, an underground cosmic ray muon experiment in the LEP ring,  LEPC99-5/P9.

[2] ALEPH Collaboration, Re: CosmoLEP Proposal, LEPC 99-7/R14.